ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

18. Teaching and Popularization of Mathematics

1991 MS Classification: 68, 94 Ozaki, Yashiro, and Sano, Kimirê, Hachinohe Insitute of Technology 88-1, Myo Ohbiraki Hachinobe, Aomori, 031-8501, Japan An Approach in Teaching by Means of the Personal Computer Software “Mathematica” This presentation introduces an improved method of teaching mathematics based upon analysis of the process by which college students require mathematical principles. By the use of rich, colorful animations, my hope is to increase the understanding of mathematics for students, by heightening their motivation. We show some examples of animation used in calculus, function, limit value, tangent line, Newton’s method, Maclaurin’s expansion, etc.